A cannon tilted upward at θ = 28.7 ∘ fires a cannonball with a speed of 92.5 m/s on level ground.

Part D

At what time after being fired does the cannonball reach it's maximum height above the ground?

Part E

What is the distance that the cannonball travels horizontally before reaching the ground again (assuming it was shot at exactly ground level with the velocity given)?
Answer #1

here,

the initial speed , u = 92.5 m/s

theta = 28.7 degree

D)

the time taken to reach the maximum height , t = u * sin(theta)/g

t = 92.5 * sin(28.7)/9.81 s

t = 4.53 s

E)

the horizontal distance travelled , R = u^2 * sin(2*theta)/g

R = 92.5^2 * sin(2*28.7) /9.81

R = 734.78 m